OpenStudy (anonymous):

Stacey and Michele both babysit in the evenings. Stacey’s hourly rate is $10 for one child plus $2 per additional child. Michele’s hourly rate is $8 for one child plus $3 per additional child. The Thompson family pays the same hourly rate for Stacey or Michele. How many children are in the Thompson family?

OpenStudy (anonymous):

M=S M= 8+3x S=10+2x 10+2x=8+3x x=?